OpenStudy (anonymous):

translate the following situation into a inequality.. the length fo the second side of a triangle is one foot less than thre the length of the first side.the lengthof the third side is two feet more than the length of the first side. if the perimeter of the triangle must be larger than 108 feet. what ar the dimensions of the triange?

OpenStudy (anonymous):

Denote the lengths of the sides as L1, L2, and L3. "the length fo the second side of a triangle is one foot less than thre the length of the first side." : L2 = L1- 1 "the lengthof the third side is two feet more than the length of the first side." : L3 = L1 + 2 "the perimeter of the triangle must be larger than 108 feet." : L1 + L2 + L3 > 108 L1 + L1 - 1 + L1 + 2 > 108 3L1 >107 L1 > 107/3, L2 >= 107/3 -1, L3 >= 107/3 + 2 Somth like this...